# Stimulated absorption of single gravitons: First light on quantum gravity

### v1

@@ Abstract
In a recent work we showed that the detection of the exchange of a single graviton between a massive quantum resonator and a gravitational wave can be achieved. Key to this ability are the experimental progress in preparing and measuring massive resonators in the quantum regime, and the correlation with independent LIGO detections of gravitational waves that induce stimulated absorption. But do stimulated single-graviton processes imply the quantization of gravity? Here we analyze this question and make a historic analogy to the early days of quantum theory. We discuss in what ways such experiments can indeed probe key features of the quantized interaction between gravity and matter, and outline five experimental tests. This capability would open the first window into experimental exploration of quantum gravity.
Based on essay written for the Gravity Research Foundation 2024 Awards for Essays on Gravitation.
@@
Introduction. The quantization of gravity implies the existence of the graviton. In practice, a graviton constitutes an indivisible quantum of energy $E=hf$ that forms gravitational waves of frequency $f$, and where $h$ is Planck’s constant. While gravitational waves have now been directly confirmed with LIGO from numerous compact binary mergers [1, 2], such waves are highly energetic and deeply in the classical regime. Nevertheless, it was recently shown by some of us that the detection of a single graviton is possible and realistic with near-future quantum technology [3]. One key element is that due to the very small scattering cross-section, of order $\sigma\sim l_{p}^{2}$ [4], where $l_{p}\approx 10^{-35}$ m is the Planck length, only very few gravitons from such a classical wave actually interact with the detector. It is therefore possible to design a resonant system that would absorb only a single graviton from a passing gravitational wave. Detecting this single event also requires the ability to prepare the quantum ground state of a massive resonator, time-continuous sensing of quantum jumps in discrete energy steps in the detector, and the cross-correlation to LIGO events to confirm absorption from the independently measured gravitational wave (it turns out neutron-star-mergers [2] are the best candidates). While challenging, these capabilities are realistically achievable, and thus single-graviton detection can be realized with kg-scale resonators [3], potentially augmented with lower-mass transducers for quantum sensing [5].
@@
In essence, such a detection would confirm the stimulated absorption of a single graviton. But how much can we learn about quantum gravity from such a stimulated event, extracted from a very energetic classical wave? This is analogous to the photoelectric effect, where stimulated absorption of photons was observed [6]. In 1905, Einstein proposed the quantization of light and applied this hypothesis to explain the photoelectric effect [7], which sparked the development of quantum theory. After much resistance, the quantum of light was fully accepted in the mid 1920ies. Today, however, it is well-known that stimulated emission such as in the photoelectric effect can be described with semi-classical optics [8]. Remarkably, a wide range of quantum phenomena can be explained through semi-classical models, in which matter is quantized but radiation is treated classically, including spontaneous emission, the Lamb shift or vacuum polarization [9]. A rigid, modern benchmark for the existence of photons is thus the measurement of quantum statistics of light, which can indicate the quantum nature of the radiation field through sub-Poissonian statistics or a vanishing intensity-intensity correlation function $g^{(2)}(0)$, because of the indivisibility of a single photon energy [10]. An even more stringent benchmark is the characterization of the quantum state of the incoming radiation, with quantum features typically attributed to a negative Wigner function rather than a negative Glauber-Sudarshan-P-function that is sufficient for non-classical statistics [11], or even the violation of Bell’s inequalities as the ultimate test of non-classicality [12]. Given these modern developments, it has been a natural question whether tests of quantum gravity can demonstrate such benchmarks. However, even tests of quantum statistics are far outside the reach of experiments [13].
@@
In this essay, instead, we argue that a more modest approach that relies only on stimulated processes can still experimentally reveal a great deal about the possible quantization of gravity. We first briefly review our results [3] that show how single gravitons can be detected through stimulated absorption, and then discuss how this can be used to infer and test aspects of quantum gravity. We draw a historic analogy to the early days of quantum mechanics [14, 15, 16], when the existence of the photon was accepted long before the development of full QED, let alone modern quantum optical tests. We carefully revisit the arguments that led to the acceptance of the photon, after tremendous resistance. We then apply these insights for tests of single-graviton exchange as we proposed, discussing what aspects of quantum gravity can be probed with such capabilities and similar experiments. Our results show that important foundational questions about the quantized interaction between matter and gravity can be probed, offering critical experimental input on quantum gravity.
@@
Single-graviton detection. Whether a graviton can be detected even in principle has been a fascinating question for a long time. The conventional answer has been that this is nearly impossible. Weinberg [4] (and later Boughn and Rothman [17]) for example, showed that atomic transitions involving gravitons would proceed roughly only once every $10^{40}$ s, while Dyson argued that LIGO would need a 37 orders of magnitude improvement in sensitivity to detect single gravitons [^1] Dyson estimates the number of gravitons in a typical gravitational wave by dividing the wave’s energy density $\rho_{E}=\frac{c^{2}}{32\pi G}h^{2}\nu^{2}$ by the energy density of a single graviton with $E=\hbar\nu$ in a cubic box of size $c/\nu$. [18], both vastly impractical. We revisited this question [3], and as summarized below, we show that inferring energy exchanges at the level of a single graviton is much more practical than previously thought. Before outlining this result, we wish to briefly comment on why neither LIGO nor individual atoms are good probes for inferring energy exchanges at the level of a single graviton. Firstly, LIGO measures displacements with very high precision, but this is not the best quantum mechanical observable to address the question of energy quantization. To infer energy exchanges in discrete amounts, one should think of a sensor that is sensitive to discrete changes in energy eigenstates, rather than continuous position. Secondly, atomic systems have traditionally been used to test quantum mechanics, but today quantum experiments are no longer confined to the atomic realm. For the interaction with gravity, the mass quadrupole is the relevant charge, thus scaling up to massive systems can be of great advantage. Finally, it is not necessary to improve the sensitivity to correspond to a single graviton strain amplitude for inferring the existence of gravitons. While low intensity measurements are crucial for many modern table top experiments and quantum information tasks, most quantum experiments in fact include bright sources. A historically important example is the photoelectric effect in which bright light shining on metal surfaces excites electrons, from which Einstein inferred the existence of a light-quantum. Dyson in fact also considered stimulated absorption for gravitons [18], but in his scenarios it remained impractical. Instead, we looked at several modifications to previous consideration, which enable single graviton detection. One key element is the use of a collective mode in a massive quantum system, as opposed to a single atom. Today such massive systems can be brought into quantum states, and quantum mechanical effects are explored on macro-scales [19]. That would amount to the use of a quantum version of classic Weber-bar detectors. But it is also necessary to combine it with appropriate quantum sensing capabilities to monitor discrete changes in energy, just as for an atom [20], as opposed to the conventional transduced position measurements. Notably there are still Weber bars in operation which look for classical gravitational wave signals, mostly at high frequencies [21], but traditional designs can operate below kHz-frequencies [22]. Finally, we circumvent the limitation of not being able to create gravitational waves by correlating the signal to independent LIGO detections: this provides heralded detection with independent confirmation of the amplitude and frequency of the incoming gravitational wave. However, because occurrence of such events is a priori unknown, a continuous quantum sensing scheme is required [23].
@@
Using this interaction Hamiltonian one can compute the absorption and emission processes. While spontaneous emission remains untenable, one can achieve stimulated emission rates as good as 1 Hz: For example with a $1800$ kg Al-cylinder, stimulated by a gravitational wave with amplitude $h=5\times 10^{-22}$. Thus in each detection event, only a few gravitons are exchanged. Importantly, we can solve the dynamics exactly, and find that an initial ground state of the massive system evolves to | $\lvert\,0\rangle\rightarrow\lvert\,\beta(t)e^{-i\omega t}\rangle$ with the coherent state amplitude
[…]
@@
The exact mass depends on the gravitational wave profile through $\chi$ as given in eq. (2), which can either be found from available data, or by an analytic approximation for binary mergers [3]. For sources like the GW170817 neutron star - neutron star merger as detected by LIGO [2], this yields an ideal mass on the order of $M_{opt}\approx 20$ kg for beryllium. Thus ground-state-cooled kg-scale detectors can be used to observe the absorption of a single graviton, if their energy levels are monitored through continuous quantum measurement. Such monitoring could also be implemented at much lower masses through transducers [5].
@@
As a simple example, for a monochromatic incident wave at angular frequency detuned from the detector’s resonant frequency by $\Delta$, and for $\Delta\ll\omega$, the excitation probability becomes $P_{0\rightarrow 1}=\frac{h^{2}ML^{2}\omega^{3}t^{2}}{4\hbar\pi^{4}}\textrm{sinc}^{2}(t\Delta/2)$. In the long-time limit and on resonance, this yields the usual Fermi’s Golden rule. This expression highlights the direct correspondence to the main features of the photoelectric effect: The probability of excitation, or the number of absorption events, increases with the intensity $h^{2}$, but the energy of transition is completely determined by the resonance condition $\Delta\approx 0$. The energy changes in discrete amounts at or near the resonant frequency. This is the well-known dependence on frequency and independence on intensity from the photoelectric effect. What is missing is the ability to measure the matter excitations through the voltage – instead one can monitor the quantum jumps directly through time-continuous quantum non-demolition measurements of energy [23] and thus deduce the exchange of a single graviton. While challenging, energy measurements in massive quantum systems, albeit at lower masses, have been demonstrated [25].
@@
Thus in summary, a single graviton detection is possible. But apart from improved technology (quantum sensing in low-noise massive resonators and ground state cooling), it relies on stimulated emission from incoming gravitational waves. It thus cannot serve as a proof of the quantization of gravity – stimulated processes can be described in the semi-classical limit. Nevertheless it can provide indirect evidence of the quantum nature of gravity. Clearly the exchange of quanta implies the change of energy in discrete amounts from the gravitational wave, thus by energy conservation assumed to hold for each individual event, one can deduce that indeed gravitons are absorbed. But it is also useful to revisit the historic arguments for the acceptance of quantum theory. This analogy is very instructive, as the experiments on gravity we envision today are in many ways technologically constrained to indirect hints of quantumness as the early tests of quantum phenomena. Thus, rather than using modern benchmarks for non-classicality, here we focus on the historic arguments and experiments around quanta of light. While it took decades for light quanta as predicted by Einstein to be accepted, no doubt remained even long before quantum optical tests were devised. Today, a direct link to quantum gravity can be made: our work showed that the exchange of quanta between matter and gravitational waves can be observed, but a probe of the quantum statistics or even the quantum state of the gravitational field seems out of reach. Revisiting the historic arguments that conclusively showed the existence of a photon, despite immense resistance in the physics community, is thus a blueprint to explore questions about quantum gravity with similarly limited experimental capabilities in the near future.
@@
Historic perspective. Historically, Planck’s black-body calculation in 1900 showed the discrete exchange of energy between light and matter, in packages of $E=hf$ [26]. Nevertheless, it was only Einstein that took the leap in 1905 to postulate that light actually consists of such discrete energy quanta, and that it is not simply limited to an anomalous interaction. Einstein realized that Planck’s law required the modification of equipartition as predicted by classical electromagnetic theory, which would have led to what is today known as the Rayleigh-Jeans black-body behavior. Instead, he focused on the experimentally verified Wien’s distribution law for high frequencies of the black-body spectrum. He realized that this law – the high-frequency limit of Planck’s distribution – can be derived as if radiation behaved thermodynamically as ideal point particles, each with energy $E=hf$. This yields the correct change in entropy as function of volume, in direct analogy to an ideal Boltzmann gas, thus suggesting a gas of particles of light. He thus made the leap from the hypothesis of a quantized interaction between light and matter, to postulating that light itself behaves as if consisting of quanta [7, 27]. Einstein applied his hypothesis of quantized interaction to the photoelectric effect. Nevertheless, despite the success of his predictions, the theory of quantized light was not accepted for over 20 years. Even after his famous measurements of Planck’s constant through Einstein’s formula for the photoelectric effect in 1916 [28], Millikan still considered Einstein’s quantized radiation hypothesis “untenable” and even “reckless” [14]. Thus, while the quantized exchange of energy was gradually accepted, the leap to a quantization of the free radiation field was a much bigger challenge.
@@
What then, beyond the photoelectric effect but long before quantum optical tests, lead to the acceptance of quantized radiation? And what lessons can we draw for the gravitational case? The developments after 1905 largely focused on the quantization of matter and the behavior of solids: Einstein realized that his quantum hypothesis, if applied to matter alone just as for light, implies a specific heat of materials that reproduced experiments at the time, but gave new predictions at low temperatures [29]. Such experiments started to be realized in the early 1910s and confirmed Einstein’s predictions, which drew attention to his quantum theory. Much of the work then focused on developing a more detailed quantum theory for matter, most notably Bohr’s atomic model. Nevertheless, a quantum theory for radiation seemed unnecessary to most, including Planck, Millikan and even Bohr himself. Even Einstein had his doubts, but by 1917 was fully convinced about the reality of light quanta. Two developments were key in the final acceptance of quantized radiation.
@@
Secondly, alternative views have become increasingly nonviable. Early attempts to explain the photoelectric effect without photons, such as Lenard’s triggering hypothesis [6], were dismissed due to experimental contradictions to such models. It was realized by Einstein and others, that if giving up energy conservation, one can explain the photoelectric effect statistically without light quanta but that this is undesirable [^2] At the first 1911 Solvay conference, Einstein remakred: “One can choose between the [quantum] structure of radiation and the negation of an absolute validity of the energy conservation law. […] We will agree that the energy principle should be retained.” [32]. This in fact remains a feature of the modern quantum optical semi-classical limit, where energy is not conserved for individual absorption or emission events [9]. In our proposal to detect single gravitons [3], we argue precisely on these grounds of energy conservation that single graviton exchange can be inferred, when single quantum jumps are recorded. However, it is possible to envision energy-non-conserving alternatives. A famous one in the context of electromagnetic radiation was put forward by Bohr, Kramers and Slater (BKS) in 1924 [33], as a means to avoid the need for photons. It attempted to keep quantized matter without quantizing the field, conserving energy and momentum on average, but violating these conservation laws for the individual exchange of quanta between field and matter. It is in this spirit similar to today’s semi-classical limit. But this theory was quickly dismissed experimentally through Bothe’s and Geiger’s coincidence measurements between scattered photons and electrons, showing event-by-event conservation of energy and momentum [34].
@@
Thus, from limited but diverse experimental input and without direct observation, it was possible to confirm the existence of quanta of light. Some lingering doubts remained, which today we might call “loopholes”. For example by Erwin Schrödinger [35, 36], who considered the possibility of a classical electromagnetic wave which would Bragg-scatter from matter-waves to explain the Compton effect, and keeping the spirit of BKS theory by abandoning conservation laws for single events. This eventually motivated modern quantum optical coincidence experiments demonstrating sub-Poissonian statistics [10, 37], anti-bunching [38] and Hong-Ou-Mandel interference [39], putting most reservations to rest. Nevertheless, it should be stressed that long before tests of quantum statistics or QED, quantization of radiation was confirmed to a high degree. While semi-classical models were able to account for some experimental features, they generally failed to capture different observations [40]. This is the benchmark we wish to apply to the gravitational case today.
Tests with single-graviton interactions. From the above historical developments and considerations, we thus motivate the following 5 tests of the quantum interaction between matter and gravitons, that are enabled by the setup we proposed:

### v2

@@ Abstract
In a recent work we showed that the detection of the exchange of a single graviton between a massive quantum resonator and a gravitational wave can be achieved. Key to this ability are the experimental progress in preparing and measuring massive resonators in the quantum regime, and the correlation with independent detections of gravitational waves (GWs) that induce stimulated absorption, such as by the Laser Interferometer Gravitational Wave Observatory (LIGO). But can stimulated single-graviton processes shed light on the quantization of gravity? Here we analyze this question and make a historic analogy to the early days of quantum theory. We discuss in what ways such experiments can indeed probe key features of the quantized interaction between gravity and matter. We outline five experimental tests utilizing stimulated emission and absorption processes of single gravitons, opening the first window into experimental exploration of the fundamentals of quantum gravity.
I Introduction
@@
The quantization of gravity implies the existence of a fundamental particle associated with the gravitational interaction — the graviton. Even though no full theory of quantum gravity has yet been found, the prediction of a graviton is a core feature in most approaches. Based on regular quantized linear gravity, a graviton is a massless boson that mediates the gravitational interaction, and which carries spin-2 and no other charge. In practice, a graviton also constitutes an indivisible quantum of energy $E=hf$ that forms gravitational waves of frequency $f$, and where $h$ is Planck’s constant. While gravitational waves have now been directly confirmed with LIGO from numerous compact binary mergers [1, 2], such waves are highly energetic and deeply in the classical regime. Nevertheless, it was recently shown by some of us that the detection of a single graviton is possible and realistic with near-future quantum technology [3]. One key element is that due to the very small scattering cross-section, of order $\sigma\sim l_{p}^{2}$ [4] for single atoms, where $l_{p}\approx 10^{-35}$ m is the Planck length, only very few gravitons from such a classical wave actually interact with the detector. It is therefore possible to design a resonant macroscopic system that would absorb only a single graviton from a passing gravitational wave — but which is still laboratory-scale. Detecting this single event also requires the ability to prepare the quantum ground state of the massive resonator, time-continuous sensing of quantum jumps in discrete energy steps in the detector, and the cross-correlation to LIGO events to confirm absorption from the independently measured gravitational wave (it turns out neutron-star-mergers [2] are the best candidates), as illustrated in Fig. 1. While challenging, these capabilities are realistically achievable, and thus single-graviton detection can be realized with kg-scale resonators [3], potentially augmented with lower-mass transducers for quantum sensing [5].
@@
Here we discuss how such graviton detection capabilities can offer first empirical input on the fundamentals of linearized quantum gravity. In essence, such a detection would confirm the stimulated absorption or emission of a single graviton. But how much can we learn about quantum gravity from such a stimulated event, extracted from a very energetic classical wave? It is helpful to draw an analogy to the early tests of quantum theory at the turn of the 20th century. Graviton detection would be analogous to the photoelectric effect, where stimulated absorption of photons was observed [6], and which is the basis of modern-day photon-detectors. In 1905, Einstein proposed the quantization of light and applied this hypothesis to explain the photoelectric effect [7], which sparked the development of quantum theory. After much resistance, the quantum of light was fully accepted in the mid 1920ies. Today, however, it is well-known that stimulated emission such as in the photoelectric effect can be described with semi-classical optics [8]. Stimulated single-quantum transitions alone do not conclusively prove that the radiation is a quantum field. Remarkably, a wide range of quantum phenomena can be explained through semi-classical models, in which matter is quantized but radiation is treated classically, including spontaneous emission, the Lamb shift, and vacuum polarization [9] (but with the drawback that conservation laws have to be abandoned). A rigid, modern benchmark for the existence of photons is thus the measurement of quantum statistics of light, which can indicate the quantum nature of the radiation field through sub-Poissonian statistics or a vanishing intensity-intensity correlation function $g^{(2)}(0)$, because of the indivisibility of a single photon energy [10]. An even more stringent benchmark is the characterization of the quantum state of the incoming radiation, with quantum features typically attributed to a negative Wigner function rather than a negative Glauber-Sudarshan-P-function that is sufficient for non-classical statistics [11], or even the violation of Bell’s inequalities as the ultimate test of non-classicality [12]. Given these modern developments, it has been a natural question whether tests of quantum gravity can demonstrate such benchmarks [^1] After the first version of this manuscript, it was shown that any potential quantum statistical features of the radiation are directly accessible in this type of detector [13, 14, 15].. Realizing some of the most stringent tests available to quantum optics today would be very challenging for the graviton case [16].
Figure 1: A single-graviton-detector consists of a massive system with quantized energy levels that can be resolved, and prepared in an initial energy eigenstate such as the ground state. The weak interaction with a passing gravitational wave can result in resonant absorption of a single graviton of energy $E=\hbar\nu$ matching the detector energy $E=\hbar\omega$. The correlation to independent detection of the passing GW indicates the stimulated single graviton process. This detector principle was outlined in Ref. [3] and may be feasible in near-future experiments.
@@
Here, instead, we argue that a more modest approach that relies only on stimulated processes can still experimentally reveal a great deal about the possible quantization of gravity. We first briefly review our results [3] that show how single gravitons can be detected through stimulated absorption, and then discuss how this can be used to infer and test aspects of quantum gravity. We draw a historic analogy to the early days of quantum mechanics [17, 18, 19], when the existence of the photon was well established and accepted by the physics community long before the development of full QED, let alone modern quantum optical tests. We then propose tests of fundamental principles of single-graviton exchange, discussing what aspects of quantum gravity can be probed with such capabilities and similar experiments. Our results show that important foundational questions about the quantized interaction between matter and gravity can be probed, offering critical experimental input on quantum gravity.
II Single-graviton detection
@@
Whether a graviton can be detected even in principle has been a fascinating question for a long time. The conventional answer has been that this is nearly impossible. Weinberg [4] (and later Boughn and Rothman [20]) for example, showed that atomic transitions involving gravitons would proceed roughly only once every $10^{40}$ s, while Dyson argued that LIGO would need a 37 orders of magnitude improvement in sensitivity to detect single gravitons (Dyson estimates the number of gravitons in a typical gravitational wave by dividing the wave’s energy density $\rho_{E}=\frac{c^{2}}{32\pi G}h^{2}\nu^{2}$ by the energy density of a single graviton with $E=\hbar\nu$ in a cubic box of size $c/\nu$) [21], both vastly impractical. We revisited this question [3], and as summarized below, we show that inferring energy exchanges at the level of a single graviton is much more practical than previously thought. Before outlining this result, we wish to briefly comment on why neither LIGO nor individual atoms are good probes for inferring energy exchanges at the level of a single graviton. Firstly, LIGO measures displacements with very high precision, but this is not the best quantum mechanical observable to address the question of energy quantization. To infer energy exchanges in discrete amounts, one should think of a sensor that is sensitive to discrete changes in energy eigenstates, rather than continuous position. Secondly, atomic systems have traditionally been used to test quantum mechanics, but today quantum experiments are no longer confined to the atomic realm. For the interaction with gravity, the mass quadrupole is the relevant charge, thus scaling up to massive systems can be of great advantage. Finally, it is not necessary to improve the sensitivity to correspond to a single graviton strain amplitude for inferring the existence of gravitons. While low intensity measurements are crucial for many modern table top experiments and quantum information tasks, most quantum experiments in fact include bright sources. A historically important example is the photoelectric effect in which bright light shining on metal surfaces excites electrons, from which Einstein inferred the existence of a light-quantum. Dyson in fact also considered stimulated absorption for gravitons [21], but in his scenarios it remained impractical. Instead, in Ref. [3] we looked at several modifications to previous consideration, which enable single graviton detection. One key element is the use of a collective mode in a massive quantum system, as opposed to a single atom. Today such massive systems can be brought into quantum states, and quantum mechanical effects are explored on macro-scales [22]. That would amount to the use of a quantum version of classic Weber-bar detectors. But it is also necessary to combine it with appropriate quantum sensing capabilities to monitor discrete changes in energy, just as for an atom [23], as opposed to the conventional transduced position measurements. Notably there are still Weber bars in operation which look for classical gravitational wave signals, mostly at high frequencies [24], but traditional designs can operate below kHz-frequencies [25]. Finally, we circumvent the limitation of not being able to create gravitational waves by correlating the signal to independent LIGO detections: this provides heralded detection with independent confirmation of the amplitude and frequency of the incoming gravitational wave. However, because occurrence of such events is a priori unknown, a continuous quantum sensing scheme is required [26].
@@
Using this interaction Hamiltonian one can compute the absorption and emission processes. While spontaneous emission remains untenable, one can achieve stimulated emission rates as good as 1 Hz: For example with a $1800$ kg Al-cylinder, stimulated by a gravitational wave with amplitude $h=5\times 10^{-22}$. Thus in each detection event, only a few gravitons are exchanged. Importantly, we can solve the dynamics exactly, and find that an initial ground state of the massive system evolves to $\lvert\,0\rangle\rightarrow\lvert\,\beta(t)e^{-i\omega t}\rangle$ with the coherent state amplitude
$|\beta(t)|=\sqrt{\frac{ML^{2}}{\pi^{4}\omega\hbar}}\,\left|\int_{0}^{t}ds\ddot{h}(s)e^{-i\omega s}\right|=\sqrt{\frac{ML^{2}}{\pi^{4}\omega\hbar}}\,\,\chi(h(t),\omega,t)\,.$ (2)
@@
The exact mass depends on the gravitational wave profile through $\chi$ as given in eq. (2), which can either be found from available data, or by an analytic approximation for binary mergers [3]. For sources like the GW170817 neutron star — neutron star merger as detected by LIGO [2], this yields an ideal mass on the order of $M_{opt}\approx 20$ kg for beryllium composition. Thus ground-state-cooled kg-scale detectors can be used to observe the absorption of a single graviton, if their energy levels are monitored through continuous quantum measurement. Such monitoring could also be implemented at much lower masses through transducers [5].
We note that noise supersession is also key, but it is feasible given the very short relevant window of cross-correlation to the classical detection events. For example, the NS-NS mergers detected at LIGO [2] last only a few seconds within the detection band, and thus false positive counts in the click detector have to be suppressed only in this time window. This can be achieved with a mechanical $Q$-factor of roughly $Q=10^{10}$ and an environment temperature of $T=1$ mK, as discussed in Ref. [3]. In contrast, the mechanical Q-factor does not pose a limitation for the detector response to the signal itself. This is because the dependence on the quality factor typically arises for the steady-state between incident wave and detector absorption — but for the aforementioned binary star mergers, this regime is never reached. Instead, the short time of the GW duration within the resonance window is of relevance (see also Appendix A), which enters in the optimal mass in eq. (4) through the GW-dependent function $\chi$.
@@
As a simple example, for a monochromatic incident wave at angular frequency detuned from the detector’s resonant frequency by $\Delta$, and for $\Delta\ll\omega$, the excitation probability becomes $P_{0\rightarrow 1}=\frac{h^{2}ML^{2}\omega^{3}t^{2}}{4\hbar\pi^{4}}\textrm{sinc}^{2}(t\Delta/2)$. In the long-time limit and on resonance, this yields the usual Fermi’s Golden rule. This expression highlights the direct correspondence to the main features of the photoelectric effect: The probability of excitation, or the number of absorption events, increases with the GW intensity $h^{2}$, but the energy of transition is completely determined by the resonance condition $\Delta\approx 0$. The energy changes in discrete amounts at or near the resonant frequency. This is the well-known dependence on frequency and independence on intensity from the photoelectric effect. What is missing is the ability to measure the matter excitations through the voltage — instead one can monitor the quantum jumps directly through time-continuous quantum non-demolition measurements of energy [26] and thus deduce the exchange of a single graviton. While challenging, energy measurements in massive quantum systems, albeit at lower masses, have been demonstrated [28].
Thus in summary, a single graviton detection is possible in the not-so-distant future, given the rapid advancements in quantum acoustics. But apart from improved technology requirements (quantum sensing in low-noise massive resonators and ground state cooling), the detection principle relies on stimulated emission from incoming gravitational waves. It thus cannot serve as a proof of the quantization of gravity — stimulated processes can be described in the semi-classical limit. This is analogous to the Rabi-limit of the interaction Hamiltonian between light and matter: the gravitational field in eq. (1) can be treated as a classical parameter, while the matter is quantized. In this limit, the same excitation probabilities are recovered. To go beyond the limit, it would be necessary to probe the quantum states of the radiation field or other quantum features, as discussed in the introduction.
@@
Nevertheless even stimulated processes can provide indirect evidence of the quantum nature of gravity, just as for the quantum nature of light in the first half of the 20th century. Clearly the exchange of quanta implies the change of energy in discrete amounts from the gravitational wave, thus assuming energy conservation to hold for each individual event, one can deduce that indeed gravitons are absorbed. But it is also useful to revisit the historic arguments for the acceptance of quantum theory. This analogy is very instructive, as the experiments on gravity we envision today are in many ways technologically constrained to indirect hints of quantumness as the early tests of quantum phenomena. Thus, rather than using modern benchmarks for non-classicality, here we focus on the historic arguments and experiments around quanta of light. This is not to suggest that the exact same reasoning would hold for gravity — but rather to revisit what level of experimental evidence was attainable to make the existence of photons likely, long before modern quantum optical tests. A direct link to quantum gravity can then be made. Our work showed that the exchange of quanta between matter and gravitational waves can be observed, but a probe of modern, stringent benchmarks of quantum behavior of gravity seems out of reach. Revisiting the historic arguments that showed early on the existence of a photon with high confidence can thus serve as a blueprint to explore questions about quantum gravity with similarly limited experimental capabilities in the near future.
III Historic perspective
@@
Historically, Planck’s black-body calculation in 1900 showed the discrete exchange of energy between light and matter, in packages of $E=hf$ [29]. Nevertheless, it was only Einstein that took the leap in 1905 to postulate that light actually consists of such discrete energy quanta, and that it is not simply limited to an anomalous interaction. Einstein realized that Planck’s law required the modification of equipartition as predicted by classical electromagnetic theory, which would have led to what is today known as the Rayleigh-Jeans black-body behavior. Instead, he focused on the experimentally verified Wien’s distribution law for high frequencies of the black-body spectrum. He realized that this law – the high-frequency limit of Planck’s distribution – can be derived as if radiation behaved thermodynamically as ideal point particles, each with energy $E=hf$. This yields the correct change in entropy as function of volume, in direct analogy to an ideal Boltzmann gas, thus suggesting a gas of particles of light. He thus made the leap from the hypothesis of a quantized interaction between light and matter, to postulating that light itself behaves as if consisting of quanta [7, 30]. Einstein applied his hypothesis of quantized interaction to the photoelectric effect. Nevertheless, despite the success of his predictions, the theory of quantized light was not accepted for close to 20 years. Even after his famous measurements of Planck’s constant through Einstein’s formula for the photoelectric effect in 1916 [31], Millikan still considered Einstein’s quantized radiation hypothesis “untenable” and even “reckless” [17]. Thus, while the quantized exchange of energy was gradually accepted, the leap to a quantization of the free radiation field was a much bigger challenge.
@@
What then, beyond the photoelectric effect but long before quantum optical tests, led to the acceptance of quantized radiation? And what lessons can we draw for the gravitational case, for which some of the modern quantum optical tests are far out of reach? The developments after 1905 largely focused on the quantization of matter and the behavior of solids: Einstein realized that his quantum hypothesis, if applied to matter alone just as for light, implies a specific heat of materials that reproduced experiments at the time, but gave new predictions at low temperatures [32]. Such experiments started to be realized in the early 1910s and confirmed Einstein’s predictions, which drew attention to his quantum theory. Much of the work then focused on developing a more detailed quantum theory for matter, most notably Bohr’s atomic model. Nevertheless, a quantum theory for radiation seemed unnecessary to most, including Planck, Millikan and even Bohr himself. Even Einstein had his doubts, but by 1917 was fully convinced about the reality of light quanta. Two developments were key in the final acceptance of quantized radiation.
@@
Secondly, alternative views have become increasingly nonviable. Early attempts to explain the photoelectric effect without photons, such as Lenard’s triggering hypothesis [6], were dismissed due to experimental contradictions to such models. It was realized by Einstein and others, that if giving up energy conservation, one can explain the photoelectric effect statistically without light quanta but that this is undesirable [^2] At the first 1911 Solvay conference, Einstein remarked: “One can choose between the [quantum] structure of radiation and the negation of an absolute validity of the energy conservation law. […] We will agree that the energy principle should be retained.” [35]. This in fact remains a feature of the modern quantum optical semi-classical limit, where energy is not conserved for individual absorption or emission events [9]. In our proposal to detect single gravitons [3], we argue precisely on these grounds of energy conservation that single graviton exchange can be inferred, when single quantum jumps are recorded. However, it is possible to envision energy-non-conserving alternatives. An early one in the context of electromagnetic radiation was put forward by Bohr, Kramers and Slater (BKS) in 1924 [36], as a means to avoid the need for photons. It attempted to keep quantized matter without quantizing the field, conserving energy and momentum on average, but violating these conservation laws for the individual exchange of quanta between field and matter. It is in this spirit similar to today’s semi-classical limit. But this theory was quickly dismissed experimentally through Bothe’s and Geiger’s coincidence measurements between scattered photons and electrons, showing event-by-event conservation of energy and momentum [37].
@@
Thus, from limited but diverse experimental input and without direct observation, it was possible to confirm the existence of quanta of light. Some lingering doubts remained, which today we might call “loopholes”. For example by Erwin Schrödinger [38, 39], who considered the possibility of a classical electromagnetic wave which would Bragg-scatter from matter-waves to explain the Compton effect, and keeping the spirit of BKS theory by abandoning conservation laws for single events. This eventually motivated modern quantum optical coincidence experiments demonstrating sub-Poissonian statistics [10, 40], anti-bunching [41] and Hong–Ou–Mandel interference [42], putting most reservations to rest. Nevertheless, it should be stressed that long before tests of quantum statistics or QED, quantization of radiation was confirmed to a high degree. While semi-classical models were able to account for some experimental features, they generally failed to capture different observations [43]. This is the benchmark we wish to apply to the gravitational case today.
IV Tests with single-graviton interactions
From the above historical developments and considerations, one can draw inspiration to design tests of quantum theory, even without the firm proof that modern quantum optical tests would offer. There is much to be explored even in the linearized limit of quantum gravity, and the single-graviton detection offers direct input on the regime of physics where gravity interacts with matter at the level of individual quanta. As briefly mentioned above, the expected quantization of Einstein’s gravity makes firm predictions on the properties of these particles: it should be a single, massless gravitational particle, propagating at speed $c$, have spin-2, and carry energy $E=hf$. For each of these particle predictions, there is a corresponding collective behavior that explains the classical domain, and so there are indirect tests and constraints on these properties from current GW observations and consistency with classical gravity [44]. However, this quantum-classical correspondence only works if one assumes the standard quantum framework. A direct test of these properties of a single graviton is conceptually different: it would give input on the standard quantum predictions and its alternatives directly at the level of individual quanta. Importantly, there are many different theories on the nature of quantum gravity, some with their own concrete predictions on the nature of gravitons (however, many approaches are still mathematical and do not give concrete physical predictions). For example, several theories on symmetry breaking with Nambu-Goldstone modes predict a finite number of massive gravitons in addition to a massless one [45, 46]. Many bi-metric theories predict two spin-2 gravitons, one massive and one massless [47, 48]. $f(R)$ models, in certain cases, predict a second particle called a scalaron to account for an additional scalar degree of freedom found in the linearized representation [49]. And entropic gravitational theories lack gravitons altogether, instead predicting that gravity is an emergent phenomenon without quantization [50, 51]. There are countless other possibilities for properties of gravitons, once the standard linearized quantum gravity framework is altered or abandoned.
Here, rather than focusing on specific speculative models, we instead present how some of the fundamental principles of linearized quantum gravity become accessible in graviton detection experiments, directly at the level of quantum theory. With some inspiration from the first experiments on quantum theory in the early 20th century, we motivate the following five tests of the quantum interaction between matter and gravitons (see also Fig. 2), that are enabled by the setup based on stimulated exchange of quanta between GWs and matter: